# exampleframelinecolor=gray!30, linewidth=0.5pt, backgroundcolor=gray!5, roundcorner=2pt, innertopmargin=8pt, innerbottommargin=8pt, innerleftmargin=8pt, innerrightmargin=8pt, skipabove=skipbelow=

### v1

@@ Appendix B Do Alignment Subspaces Encode Safety?
Appendix C Can Harmful Subspaces Be Removed?
Table 4 presents supplementary results that further substantiate the findings discussed in Section 4.
@@ Appendix C Can Harmful Subspaces Be Removed?
Model Method Utility ($\uparrow$) emphSVD Fractions Harmful Score ($\downarrow$) emphSVD Fractions
@@
Appendix D Do Safety Subspaces Exist in Representation Space?
@@
Appendix E Experimental Details
We implemented all experiments using PyTorch [38] and the HuggingFace Transformers library [56]. We ran all experiments on a single NVIDIA A6000 GPU (48 GB). To save memory, all base models are initalized in torch.bfloat16 precision. All models are trained using the AdamW optimizer [34]. Detailed hyperparameter configurations for full fine-tuning of each model are presented in Table 5.
Optimizer AdamW
Batch size 1
Max. Seq. Len 512
Grad Acc. Steps 32
Epochs 1
Learning Rate $1\times 10^{-5}$
LR Scheduler Cosine
Warmup Ratio 0.02
Appendix F Dataset Details
We use the MetaMathQA dataset [63] for fine-tuning, which reformulates existing math problems from alternative perspectives without introducing new content. To evaluate performance, we rely on the GSM8K benchmark [8], a dataset of elementary-level math questions that require multi-step reasoning. Models are assessed based solely on the correctness of the final numerical answer. For our activation-based analysis, we sample prompts from the MATH dataset [17], which contains challenging, competition-style arithmetic problems.
BeaverTails [26] is a valuable dataset for studying safety by independently annotating question–answer pairs for both helpfulness and harmlessness. We use the training set to fine-tune models in both harmful and contaminated settings, and draw prompts from the test split for our activation-based experiments.
AdvBench [68] consists of 500 prompts designed to elicit a wide range of harmful behaviors, including profanity, threats, misinformation, discrimination, cybercrime, and other forms of dangerous or illegal content framed as instructions. We use this benchmark to quantify model harmfulness: higher success in responding to these prompts indicates greater unsafe behavior.
ToxiGen [14] is a large-scale dataset composed of both toxic and non-toxic statements. We use a subset of its prompts to analyze model activations in response to harmful content.

### v2

@@ Appendix C Do Alignment Subspaces Encode Safety?
Llama-3.2 1B Base $0.026$ $4.13$
Safety-Tuned $0.032$ $3.41$
Fine-Tuned $0.075$ $3.79$
Top-K $0.026$ $0.026$ $0.037$ $0.041$ $0.039$ $3.47$ $3.55$ $3.63$ $3.71$ $3.69$
Random-K $0.031$ $0.026$ $0.033$ $0.037$ $0.042$ $3.46$ $3.46$ $3.54$ $3.66$ $3.69$
Random $0.031$ $0.026$ $0.030$ $0.038$ $0.039$ $3.44$ $3.48$ $3.54$ $3.61$ $3.66$
@@
Llama-2 7B Base $0.053$ $4.27$
Safety-Tuned $0.024$ $3.54$
Fine-Tuned $0.151$ $3.84$
Top-K $0.030$ $0.042$ $0.058$ $0.082$ $0.122$ $3.54$ $3.67$ $3.70$ $3.78$ $3.79$
Random-K $0.026$ $0.029$ $0.043$ $0.067$ $0.117$ $3.55$ $3.59$ $3.66$ $3.75$ $3.79$
Random $0.026$ $0.033$ $0.042$ $0.064$ $0.112$ $3.55$ $3.58$ $3.68$ $3.76$ $3.79$
Appendix D Can Harmful Subspaces Be Removed?
Figure 7 and Table 5 presents supplementary results that further substantiate the findings discussed in Section 4.
@@ Appendix D Can Harmful Subspaces Be Removed?
Llama-3.2 1B Base $0.026$ $4.13$
Safety-Tuned $0.032$ $3.41$
Fine-Tuned $0.068$ $3.65$
Top-K $0.027$ $0.026$ $0.033$ $0.048$ $0.039$ $3.42$ $3.48$ $3.59$ $3.56$ $3.62$
Random-K $0.030$ $0.026$ $0.033$ $0.042$ $0.040$ $3.43$ $3.44$ $3.40$ $3.46$ $3.60$
Random $0.032$ $0.026$ $0.032$ $0.039$ $0.039$ $3.42$ $3.49$ $3.44$ $3.53$ $3.59$
@@
Llama-2 7B Base $0.053$ $4.27$
Safety-Tuned $0.024$ $3.54$
Fine-Tuned $0.168$ $3.89$
Top-K $0.029$ $0.036$ $0.055$ $0.082$ $0.127$ $3.49$ $3.69$ $3.76$ $3.83$ $3.85$
Random-K $0.026$ $0.032$ $0.042$ $0.062$ $0.128$ $3.59$ $3.58$ $3.75$ $3.79$ $3.89$
Random $0.026$ $0.036$ $0.041$ $0.065$ $0.122$ $3.57$ $3.64$ $3.74$ $3.78$ $3.89$
Appendix E Are Safety Weight Subspaces Distinct?
To supplement the analysis in Section 5, we report extended results in Figure 8.
Figure 8: Mode Subspace Overlap (MSO) at the 70-, 80-, and 85- percentile layers for pairwise comparisons of the dominant subspaces from Harmful fine-tuned (H), Safety-Tuned (S), and Base (B) models.
Appendix F Do Safety Subspaces Exist in Activation Space?
@@
Appendix G Dataset Details
We use the MetaMathQA dataset (67) for fine-tuning, which reformulates existing math problems from alternative perspectives without introducing new content. To evaluate performance, we rely on the GSM8K benchmark (9), a dataset of elementary-level math questions that require multi-step reasoning. Models are assessed based solely on the correctness of the final numerical answer. For our activation-based analysis, we sample prompts from the MATH dataset (18), which contains challenging, competition-style arithmetic problems.
BeaverTails (28) is a valuable dataset for studying safety by independently annotating question–answer pairs for both helpfulness and harmlessness. We use the training set to fine-tune models in both harmful and contaminated settings, and draw prompts from the test split for our activation-based experiments.
AdvBench (72) consists of 500 prompts designed to elicit a wide range of harmful behaviors, including profanity, threats, misinformation, discrimination, cybercrime, and other forms of dangerous or illegal content framed as instructions. We use this benchmark to quantify model harmfulness: higher success in responding to these prompts indicates greater unsafe behavior.
ToxiGen (15) is a large-scale dataset composed of both toxic and non-toxic statements. We use a subset of its prompts to analyze model activations in response to harmful content.
Appendix H Limitations
Our study focuses on linear subspaces, providing a principled first step toward understanding the geometric structure of safety alignment. While we do not explore non-linear representations, our framework could be extended in future work to capture richer geometric phenomena. Our experiments are restricted to open-weight models with publicly available base and aligned variants. These models provide a controlled and interpretable setting, though extending to production-level closed-source models remains an important direction for future work.
Appendix I Use of Large Language Models
Our use of LLMs is restricted to light writing assistance, including grammar polishing and enhancing clarity.